OpenStudy (anonymous):

What is the equation of the quadratic graph with a focus of (4, 0) and a directrix of y = 10?

OpenStudy (tkhunny):

This is enough information to immediately find the Vertex. Draw the given information and then just write in the Vertex. Go!

OpenStudy (anonymous):

Question 19 (Multiple Choice Worth 1 points) (03.09) What is the equation of the quadratic graph with a focus of (4, 0) and a directrix of y = 10? f(x) = −one twentieth (x − 4)2 + 5 f(x) = −one twentieth (x + 4)2 + 5 f(x) = one twentieth (x − 4)2 + 5 f(x) = one twentieth (x + 4)2 + 5